OpenStudy (anonymous):

Find the sum of each arithmetic series. the first 100 odd natural numbers.

OpenStudy (anonymous):

@amistre64 @zzr0ck3r @radar

OpenStudy (zzr0ck3r):

\[\sum_{n=1}^{100}2n-1=\sum_{n=1}^{100}2n-\sum_{n=1}^{100}1=2\sum_{n=1}^{100}n-\sum_{n=1}^{100}1\]

OpenStudy (zzr0ck3r):

\[2\frac{100*101}{2}-100\]

OpenStudy (zzr0ck3r):

100*100 = 10000

OpenStudy (zzr0ck3r):

I stole @amistre64 method:)

OpenStudy (zzr0ck3r):

1+3+5+....+99 199+197+.....101 add them top to bottom 200*50 = 10000

OpenStudy (anonymous):

Thanks

OpenStudy (zzr0ck3r):

np